# On Early Detection of Hallucinations in Factual Question Answering

### v2

@@ 5.2. Hallucination Classifiers
Table 2 shows the test set Area Under the Receiver Operator Curve (AUROC) of classifiers described in Section 3.3 in detecting hallucinations. We opt for AUROC instead of binary classification accuracy because of high class imbalance (Table 1).
Results—in Table 2 for larger model variants and in Table 3 for smaller model variants—show the input attribution does slightly better than random chance on the TriviaQA dataset, but no better than random on the subject specific datasets. Softmax, on the other hand, consistently does better than random for everything except the place of birth task.
Self-attention scores and fully-connected activations outperform both saliency and Softmax for classifying hallucinated responses. Interestingly, these results hold even when $96\%$ of generated responses start with the same new line formatting character, indicating that the classifier is not simply learning tokens that correlate with hallucination, but that even within the same token, the model internal state between hallucinated and non-hallucinated results differs.
@@
Combining over datasets and artifacts. A classifier trained on all four datasets combined (labeled as “Combined” in Tables 2 and 3) tended to perform slightly better than the individual datasets. However, a classifier trained on all model artifacts combined (Softmax, IG, self-attention and fully-connected activations) showed no improvement beyond models trained on each artifact individually (Table 4). There could be several reasons for this: the input dimensionality of the classifier combining all the artifacts is much higher than the individual classifiers. The combined classifier is architecturally more complex as it consists of both dense and recurrent units. Training this mixed architecture might require special considerations. We leave the detailed analysis to a follow up work.
Figure 5. [FAL-40B on TriviaQA dataset] AUROC of the hallucination detectors using self-attention and fully-connected activations at different layers. The performance is better at later layers.
@@
TriviaQA 0.751 0.639 0.652
Capitals 0.692 0.554 0.624
Founders 0.664 0.527 0.626
Birth Place 0.533 0.589 0.683
Combined 0.577 0.578 0.601
@@
Table 4. Test AUROC of the classifier trained on all four model artifacts at once. The performance is not better than the case when individual artifacts are considered (Tables 2 and 3).
Effect of Transformer layer choice. The results so far were based on using the self-attention and fully-connected activations from the last Transformer layer. We also investigate how the performance would change as a results of a change in layer number. Results in Figure 5 show that the performance of our classifier improves with depth of the model layers. Early layers do only slightly better than random chance, while later layers shows significant improvement beyond random chance.
6. Conclusion
@@
In this paper, we address the problematic behavior where large language models hallucinate incorrect facts. By extracting different generation artifacts such as self-attention scores at the start of the response generation, we build classifiers capable of identifying hallucinations even before they occur. The classifiers are able to detect hallucinations even when the initial generated token is a formatting character. Further, we show that the ability to identify hallucinations persists across topics like capitals, founders and places of birth; and in a broader trivia knowledge context. We recommend attaching classifiers of this type to deployed LLMs as a method of flagging potentially incorrect information before it reaches the user. We also recommend model developers pay closer attention to formatting tokens like new-line characters as activations of these commonly overlooked tokens carry additional information about the model behavior. Experimenting with more elaborate search settings, e.g. those using retrieval augmented generation (lewis2020retrieval, 18), a wider range of datasets, and searching over better architectures of hallucination detectors are interesting avenues for future work.

### v3

@@ 5.2. Hallucination Classifiers
Tables 2 and 3 shows the test AUROC when detecting hallucinations. We opt for AUROC instead of binary classification accuracy because of high class imbalance (Table 1). We include the results for binary classification accuracy in Appendix A.
Results show that the IG attribution does slightly better than random chance on the TriviaQA dataset, but no better than random on the subject specific datasets. Softmax consistently does better than random for all tasks. Self-attention scores and fully-connected activations outperform both IG and Softmax. Interestingly, these results hold even though $96\%$ of generated responses start with the newline character, indicating that the classifier is not simply learning tokens that correlate with hallucination, but that even with the same token, the model internal state between hallucinated and non-hallucinated results differs.
@@
A classifier trained on all four datasets combined (labeled as “Combined” in Tables 2 and 3) tended to perform slightly better than the individual datasets. However, a classifier trained on all model artifacts combined (Softmax, IG, self-attention and fully-connected activations) showed no improvement beyond models trained on each artifact individually (Table 4). There could be several reasons for this: the input dimensionality of this combined classifier is much higher than the individual classifiers and it is architecturally more complex as it consists of both dense and recurrent units. Training this mixed architecture might require special considerations. We leave the detailed analysis to a follow up work.
5.3. Hyperparameters and baseline comparison
We take a closer look at various hyperparameters and also conduct a performance comparison with an existing method.
Effect of Transformer layer choice. The results in §5.2 were based on using the self-attention and fully-connected activations from the last Transformer layer. We also investigate how the performance would change as a results of a change in layer number. Results in Figure 5 show that the performance of our classifier improves with depth of the model layers. Early layers do only slightly better than random chance, while later layers shows significant improvement beyond random chance.
Effect of classifier hyperparameters. Given the simplicity of our hallucination detection classifiers, the only hyperparameters available were batch size, learning rate, and weight decay. We hand-selected the defaults in §4.3. We also tried (TriviaQA with FAL-40B) a range of different options for these hyperparameters where we varied batch size from $4$ to $256$, learning rate from $10^{-6}$ to $10^{-2}$ and weight decay from $10^{-4}$ to $10^{-1}$. We found that the smallest batch size and largest learning rate performed worse. But smaller adjustments (batch size $128$ vs $256$, learning rate $1\times 10^{-4}$ v.s. $2\times 10^{-4}$) made no discernible difference in performance.
We also tested larger models for both our GRU and MLP architectures. On the GRU, we tested from $4$ to $12$ gated recurrent layers. On our MLP, we tested up to $8$ layers, with widths from $32$ to $256$. In both cases, we found that larger models provided no benefit in either AUROC or accuracy. For example, for the GRU model FAL-40B with TriviaQA, expanding to $12$ recurrent layers yielded an AUC of $0.46$, slightly worse than the $0.48$ reported with $4$ layers reported in Table 2. For MLP, expanding to $8$ layers yielded $0.72$, only slightly better than the $0.71$ we reported with the single layer model.
A systematic study of these hyperparameter choices, and understanding the effect of more advanced detector architectures (e.g., Transformers, LSTMs) on the detection performance is a promising future direction.
@@ 5.3. Hyperparameters and baseline comparison
TriviaQA 0.75 0.64 0.65
Capitals 0.69 0.55 0.62
Founders 0.66 0.53 0.63
Birth Place 0.53 0.59 0.68
Combined 0.58 0.58 0.60
@@
Table 4. Test AUROC of the classifier trained on all four artifacts at once. The performance is not better than the case when individual artifacts are considered (Tables 2 and 3).
Comparison with SelfCheckGPT. We compare the performance of our method to SelfCheckGPT (manakul2023selfcheckgpt). Results are shown in Table 5 for larger model variants and Table 6 for smaller model variants. The performance is generally worse than our classifiers. We use two variants of SelfCheckGPT: BERTScore and n-gram. We use $20$ generations per input and set the temperature to $0.1$. Preliminary analysis in Table 7 shows that setting the temperature to $1.0$ as suggested in the paper (manakul2023selfcheckgpt) does not lead to better performance.
LAM-13B OPT-30B FAL-40B
TriviaQA 0.56 0.58 0.50
Capitals 0.53 0.51 0.56
Founders 0.55 0.49 0.57
Birth Place 0.51 0.54 0.48
(a) BERTScore
LAM-13B OPT-30B FAL-40B
TriviaQA 0.54 0.53 0.54
Capitals 0.54 0.55 0.54
Founders 0.58 0.56 0.57
Birth Place 0.51 0.49 0.52
(b) n-gram
Table 5. [Larger model variants] AUROC of SelfCheckGPT based on BERTScore and n-gram scores.
LAM-7B OPT-6.7B FAL-7B
TriviaQA 0.55 0.56 0.55
Capitals 0.52 0.50 0.49
Founders 0.58 0.58 0.54
Birth Place 0.52 0.51 0.56
(a) BERTScore
LAM-7B OPT-6.7B FAL-7B
TriviaQA 0.53 0.58 0.50
Capitals 0.53 0.48 0.49
Founders 0.53 0.43 0.46
Birth Place 0.57 0.58 0.55
(b) n-gram
Table 6. [Smaller model variants] AUROC of SelfCheckGPT based on BERTScore and n-gram scores.
BERTScore n-gram
TriviaQA 0.52 0.55
Capitals 0.47 0.50
Founders 0.44 0.51
Birth Place 0.55 0.48
Table 7. [FAL-7B ] AUROC of SelfCheckGPT with $T=1$.
6. Conclusion, Discussion & Limitations
@@
We address the problematic behavior where large language models hallucinate incorrect facts. By extracting different generation artifacts such as Integrated Gradients feature attribution scores and self-attention scores at various Transformer layers we build simple classifiers for detecting hallucinations. The detection can already be performed at the start of the response generation, that is, factual hallucinations can often be detected even before they occur. We show that the ability to identify hallucinations persists across different topics like Capitals, Founders and Birth Place; and in a broader trivia knowledge context. Attaching classifiers of this type to deployed LLMs can serve as an effective method of flagging potentially incorrect information before it reaches the end-users or downstream applications.
[…]
We only considered factual hallucinations in this work. It would be interesting to study if the proposed method also extends to other forms of hallucinations (huang2023survey; ji2023survey). We also considered a rather simple setup where the model output is supposed to contains a single fact only. Extending our method to probe multiple hallucinations (e.g., multiple facts in a biography like place of birth, date of birth, alma mater) in a single generation is also an important follow up direction.
Our method cannot be used with “blackbox” models that do not reveal model internals and gradients to the users. Text-only methods like SelfCheckGPT are well-suited for such situations but compensate for the lack of internal access by querying the model several times. In our case, all except one artifact can be computed “for free” during the forward pass. Nonetheless, our method can still be deployed by the model providers internally. Experimenting with more elaborate retrieval settings (e.g., those using retrieval augmented generation (lewis2020retrieval), instruction tuned models (ouyang2022training) and in-context examples (wu2024towards)), a wider range of datasets and model types is also a promising avenue for future work.
Finally, the goal of this paper was not to extract the maximum possible detection performance, but to test if certain artifacts can be promising in detecting hallucinations. For this reason, we use relatively simple detection architectures. We leave a more in-depth analysis of detection architectures to a future study.
The code repository for the paper is available at: https://github.com/amazon-science/llm-hallucinations-factual-qa.
[…]
@@ Appendix A Additional Results
Recall that due to the imbalance between hallucinations vs. non-hallucinations (Section 4.4), we reported AUROC in the main experiments (Section 5.2). Tables 8 and 9 show the binary classification accuracy for larger and smaller model variants.
LAM-13B OPT-30B FAL-40B
TriviaQA 0.64 0.67 0.66
Capitals 0.68 0.65 0.63
Founders 0.70 0.69 0.68
Birth Place 0.85 0.87 0.89
(a) IG attributions
LAM-13B OPT-30B FAL-40B
TriviaQA 0.65 0.63 0.68
Capitals 0.70 0.71 0.63
Founders 0.73 0.68 0.67
Birth Place 0.86 0.91 0.92
(b) Softmax probabilities
LAM-13B OPT-30B FAL-40B
TriviaQA 0.69 0.65 0.70
Capitals 0.70 0.71 0.66
Founders 0.72 0.74 0.68
Birth Place 0.89 0.91 0.93
(c) Self-attention scores
LAM-13B OPT-30B FAL-40B
TriviaQA 0.67 0.65 0.69
Capitals 0.69 0.70 0.65
Founders 0.73 0.74 0.68
Birth Place 0.88 0.89 0.94
(d) Fully-connected activations
Table 8. [Larger model variants] Test accuracy of binary classifiers in detecting hallucinations.
LAM-7B OPT-6.7B FAL-7B
TriviaQA 0.63 0.68 0.63
Capitals 0.70 0.64 0.67
Founders 0.69 0.70 0.68
Birth Place 0.83 0.88 0.89
(a) IG attributions
LAM-7B OPT-6.7B FAL-7B
TriviaQA 0.66 0.63 0.66
Capitals 0.67 0.68 0.64
Founders 0.71 0.63 0.68
Birth Place 0.87 0.90 0.89
(b) Softmax probabilities
LAM-7B OPT-6.7B FAL-7B
TriviaQA 0.65 0.62 0.72
Capitals 0.70 0.72 0.65
Founders 0.69 0.71 0.69
Birth Place 0.91 0.90 0.91
(c) Self-attention scores
LAM-7B OPT-6.7B FAL-7B
TriviaQA 0.66 0.68 0.66
Capitals 0.70 0.69 0.65
Founders 0.71 0.72 0.70
Birth Place 0.87 0.88 0.93
(d) Fully-connected activations
Table 9. [Smaller model variants] Test accuracy of binary classifiers in detecting hallucinations.